. Two charges and B of magnitudes +8.0*10^6C each,are kept at points (2,1) and (2,7) respectively in a Cartesian plane measured in meters .if the charges are in vacuum , calculate the force of A on B*
1
Expert's answer
2020-05-18T19:56:16-0400

Coulomb’s law describes the electrostatic force (or electric force) between two charged particles. If the particles have charges "q_1" and "q_2," are separated by distance "r," and are at rest (or moving only slowly) relative to each other, then the magnitude of the force acting on each due to the other is given by


"\bar{F}={1\over 4\pi \varepsilon_0\varepsilon}\cdot{q_1q_2\over r^3}\bar{r}"

Given

"q_1=q_2=+8.0\times10^{-6} C,"

"(x_A, y_A)=(2, 1), (x_B, y_B)=(2, 7)"

"\varepsilon=1"

"\dfrac{1}{4\pi \varepsilon_0}=9\times10^9\ \dfrac{N\cdot m^2}{C^2}"

Then


"\bar{r}=(x_B-x_A, y_B-y_A)=(2-2, 7-1)=(0,6)"

"t=|\bar{r}|=\sqrt{(0)^2+(6)^2}=6"

"|\bar{F}|=9\times10^9\ {N\cdot m^2\over C^2}\cdot{(8.0\times10^{-6} C)^2\over (6 m)^2}=1.6\times10^{-2} \ C"

The magnitude of the electrostatic force of repulsion is equal to "1.6\times10^{-2} \ C."